Journal of the Association for Computing Machinery, Vol. 43, No. 1, January 1996, pp. 1–19.

Local Management of a Global Resource in a Communication Network

YEHUDA AFEK

BARUCH AWERBUCH

SERGE PLOTKIN

AND

MICHAEL SAKS

Abstract. This paper introduces a new distributed data object called Resource Controller that provides an abstraction for managing the consumption of a global resource in a distributed system. Examples of resources that may be managed by such an object include number of messages sent, number of nodes participating in the protocol, and total CPU time consumed.

The Resource Controller object is accessed through a procedure that can be invoked at any node in the network. Before consuming a unit of resource at some node, the controlled algorithm should invoke the procedure at this node, requesting a permit to consume a unit of the resource. The procedure returns either a permit or a rejection.

The key characteristics of the Resource Controller object are the constraints that it imposes on the global resource consumption. An (M, W)-Controller guarantees that the total number of permits granted is at most M; it also ensures that, if a request is rejected, then at least M − W permits are eventually granted, even if no more requests are made after the rejected one.

In this paper, we describe several message and space-efficient implementations of the Resource Controller object. In particular, we present an (M, W)-Controller whose message complexity is O(n log^{2} n log(M/(W + 1))) where n is the total number of nodes. This is in contrast to the O(nM) message complexity of a fully centralized controller which maintains a global counter of the number of granted permits at some distinguished node and relays all the requests to that node.

Categories and Subject Descriptors. C.2.3 [Computer-Communication Networks]: Network Operations; C.2.4 [Computer-Communication Networks]: Distributed Systems; F.2.2 [Analysis of Algorithms and Problem Complexity]: Nonnumerical Algorithms and Problems; G.2.2 [Discrete Mathematics]: Graph Theory

General Terms. Algorithms, Theory

Additional Key words and Phrases. Diffusing computations, distributed computation, distributed network management, resource management

Copyright 1996 by ACM, Inc. Permission to make digital or hard copies of part or all of this work for personal or classroom use is granted without fee provided that copies are not made or distributed for profit or direct commercial advantage and that copies bear this notice and the full citation on the first page. Copyrights for components of this work owned by others than ACM must be honored. Abstracting with credit is permitted. To copy otherwise, to republish, to post on servers, or to redistribute to lists, requires prior specific permission and/or a fee. Request permissions from Publications Dept, ACM Inc., fax +1 (212) 869-0481, or permissions@acm.org.